# Part I: Tricks or Traps? A Deep Dive into RL for LLM Reasoning

### v1

@@
Abstract Reinforcement learning for LLM reasoning has rapidly emerged as a prominent research area, marked by a significant surge in related studies on both algorithmic innovations and practical applications. Despite this progress, several critical challenges remain, including the absence of standardized guidelines for employing RL techniques and a fragmented understanding of their underlying mechanisms. Additionally, inconsistent experimental settings, variations in training data, and differences in model initialization have led to conflicting conclusions, obscuring the key characteristics of these techniques and creating confusion among practitioners when selecting appropriate techniques. This paper systematically reviews widely adopted RL techniques through rigorous reproductions and isolated evaluations within a unified open-source framework. We analyze the internal mechanisms, applicable scenarios, and core principles of each technique through fine-grained experiments, including datasets of varying difficulty, model sizes, and architectures. Based on these insights, we present clear guidelines for selecting RL techniques tailored to specific setups, and provide a reliable roadmap for practitioners navigating the RL for the LLM domain. Finally, we reveal that a minimalist combination of two techniques can unlock the learning capability of critic-free policies using vanilla PPO loss. The results demonstrate that our simple combination consistently improves performance, surpassing strategies like GRPO and DAPO.
@@ 1 Introduction
Recent breakthroughs in large language models (LLMs) such as OpenAI o1 (Wu et al., 2024) and DeepSeek R1 (Shao et al., 2024) have positioned reinforcement learning (RL) as a key driver in unlocking advanced reasoning capabilities within LLMs. This is particularly evident in challenging reasoning tasks like mathematical reasoning (He et al., 2025a) and code generation (Zhuo et al., 2025), where RL has demonstrated the potential to elevate LLM performance beyond what pre-training alone can achieve. Such an emerging trend has sparked widespread interest within the research community in the direction of "RL for LLM" (or RL4LLM). In 2025, RL4LLM experienced a surge in research activity, leading to hundreds of publications across arXiv and major conferences, covering a wide range of topics from algorithmic innovation to practical engineering solutions.
@@
However, this rapid progress is shadowed by the lack of usage guidelines for existing RL techniques or tricks (Huang et al., 2024a) as well as the absence of in-depth analysis of their underlying mechanisms. Specifically, these limitations can manifest as confusion among practitioners in choosing RL tricks, as different papers provide different solutions to the same problem. For instance, GRPO (Shao et al., 2024) advocates for group-level normalization to enhance policy stability, whereas REINFORCE++ (Hu et al., 2025) argues that batch-level normalization works better. Moreover, GRPO incorporates variance in normalization, yet Dr. GRPO (Liu et al., 2025a) explicitly recommends removing variance normalization to prevent bias. Another example: GRPO (Shao et al., 2024) has achieved a breakthrough in performance through the strategy of using response-level loss calculation, while DAPO (Yu et al., 2025) has instead adopted token-level loss calculation. Such contradictory and chaotic phenomena underscore the fragmented understanding and inconsistent recommendations within the RL4LLM community. One possible reason for the above phenomenon is that the experimental settings, training data, and initialization of the existing work all have significant differences, which may also cause deviations in the summary of the conclusions.
@@
Apart from the confusion caused by the intrinsic differences of similar techniques, the numerous and seemingly orthogonal techniques, including Normalization, Clip, and Overlong Filtering, also increase the complexity of algorithm application in practice. Practitioners face non-trivial challenges in selecting an appropriate combination from a wide range of techniques to unlock the learning capacity of LLMs in specific scenarios. These ambiguities have naturally triggered a key requirement of practitioners:
@@
What scenarios are the existing techniques respectively suitable for? Is there a simple and generalized combination that can be used to enhance policy optimization?
@@
Aligned with classic RL mechanism analysis methodologies (Andrychowicz et al., 2020; Engstrom et al., 2020; Huang et al., 2024a), we systematically review the widely used RL techniques by reproducing them and independently evaluating the actual impact of each technique, based on the same open-source infrastructure framework and policy models. To comprehensively cover practical scenarios, we design extensive experimental settings incorporating datasets of varying difficulty levels, diverse model sizes, and distinct model types. Furthermore, we conduct an in-depth analysis of their theoretical foundations, implementation details, and applicable scenarios as demons. The intuitive contribution is illustrated in Figure 1. Specifically, ❶ our empirical results reveal that most RL techniques exhibit obvious preferences and sensitivities to the experimental setup, e.g., model type, data distribution, reward mechanism and hyperparameter. ❷ Based on the isolated analysis under our setup, we demonstrate that employing only two techniques, i.e., advantage normalization (group-level mean, batch-level std) and token-level loss aggregation, can unlock the learning capability of critic-free policies using vanilla PPO loss, surpassing mainstream RL4LLM algorithms incorporating redundant components. Our core contributions are selected as:
@@
1. Group-level normalization shows robust efficiency under each reward setting. Batch-level normalization provides more stable improvement under large scale reward setting. (§4.1.1) 2. Group-level mean and batch-level standard deviation enable further robust normalization. (§4.1.3) 3. Clip Higher prefers promoting high-quality exploration for aligned models. (§4.2.1) 4. There appears to be a “scaling law” between the performance and the upper bound of the clipping on the small-sized model. (§4.2.3) 5. Compared to sequence-level loss aggregation, token-level aggregation is effective on base models, while showing limited improvement on aligned models. (§4.3.1) 6. Overlong filtering enhances accuracy and clarity for short-to-medium reasoning tasks but provides limited benefits for long-tail reasoning. (§4.4.1) 7. Two techniques may unlock learning capacity in critic-free policies based on vanilla PPO loss. (§5)
@@ 2.2 Group Relative Policy Optimization (GRPO)
The effectiveness of the above normalization method can be understood from the perspective of reward shaping. By emphasizing the differences among candidate outputs for the same prompt, it effectively preserves the reliability of the gradient signal, even in sparse reward settings (Hu et al., 2020). Instead of adding KL penalty in the reward, GRPO directly regularizes by directly adding the KL divergence between the trained policy and the reference policy to the loss. The overall surrogate objective is:
@@ 2.3 Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO)
Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025) is a recent RL method designed to address the unique challenges in LLM reasoning.
For each question $q$ with gold answer $a$, DAPO samples a group of $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy, computes their rewards, and maximizes the following surrogate objective:
@@ 2.4 Reinforcement Learning Techniques
A variety of practical techniques have been introduced to stabilize optimization, reduce variance, and accelerate convergence of LLM on the reasoning task. Drawing from prior research and practical implementations, we categorize widely used techniques as follows.
@@
Clipping controls excessive updates in policy optimization and can be applied to different quantities, such as rewards, advantages, or ratios. Furthermore, the Clip Ratio Higher (Yu et al., 2025) method relaxes the upper bound in PPO’s ratio clipping to better preserve exploration.
@@
Normalization of rewards or advantages helps stabilize gradient magnitudes. Representative approaches include: Batch-level Reward Normalization (Hu et al., 2025), Group-level Reward Normalization (Shao et al., 2024; Ahmadian et al., 2024), and Reward Shift without Standard Deviation (Liu et al., 2025a), which removes the standard deviation term to avoid the difficulty bias.
@@
Filtering out uninformative or undesirable samples prior to gradient computation. Examples include: Overlong Filtering (Yu et al., 2025) to remove responses exceeding predefined length limits; Error Max Clip Mask and Right Min Clip Mask to filter overly incorrect or trivially correct samples; and Difficulty Mask (Yu et al., 2025; Zhang et al., 2025; Chu et al., 2025) to exclude samples outside a targeted difficulty range.
@@
The formulation of loss aggregation determines the relative weight each token contributes to the overall objective. Common approaches include: Token-level Loss computes per-token advantages to reduce length bias, while Sequence-level Loss aggregates at the sequence level.
@@
The above categories summarize the most prevalent improvement strategies for RL in LLM reasoning. In this work, we focus on four key aspects: Normalization, Clipping, Masking, and Loss Aggregation, and conduct in-depth analyses of their mechanisms and practical utility.
@@ 3.1 Experimental Setup
We utilize the open-sourced ROLL framework [^1] Open source RL framework: https://github.com/alibaba/ROLL (Wang et al., 2025), an efficient and scalable platform specifically designed for reinforcement learning optimization in LLMs, to conduct all experiments. Besides, we adopt PPO loss (Schulman et al., 2017), with advantage values computed using the REINFORCE algorithm (Sutton et al., 1999) as the unified and naive RL baseline. To ensure that the batch size for global sampling is consistent with existing research, i.e., $1024$, we set the rollout batch size to $128$ and sample $8$ responses for each prompt, with a maximum response length of $8192$ tokens. The learning rate is set to 1e61e-6. For text generation, we use a top_p value of $0.99$, a top_k value of $100$, and a temperature of $0.99$.
@@
To comprehensively evaluate reinforcement learning (RL) techniques across parameter scales, our experiments cover two model sizes: Qwen3-4B and Qwen3-8B. For each model size, we include both non-aligned pre-trained versions (Qwen3-4B-Base and Qwen3-8B-Base) and aligned versions, allowing us to assess RL gains from various starting points [^2] Checkpoint links: https://huggingface.co/Qwen/Qwen3-4B; https://huggingface.co/Qwen/Qwen3-8B; https://huggingface.co/Qwen/Qwen3-4B-Base;https://huggingface.co/Qwen/Qwen3-8B-Base.
@@
To ensure reproducibility and fairness, we exclusively use open-source datasets for training, including SimpleRL-Zoo-Data (Zeng et al., 2025) and Deepmath (He et al., 2025a). To comprehensively examine how problem difficulty affects the RL technique’s performance, we randomly sample from the datasets while removing an excessive proportion of examples whose ground-truth label is simply “True” or “False”. Because we identify the ostensible positive phenomenon wherein models produce correct binary answers from erroneous reasoning chains, introducing noisy supervision that compromises training quality (please refer to Appendix B.2 for case studies). Figure 2 visualizes the difficulty across the training dataset assessed by GPT-4o (Hurst et al., 2024).
@@
All the experiments are conducted on six math datasets. i.e., MATH-500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024), MinervaMath (Lewkowycz et al., 2022), and subsets of standardized examinations (AIME24-25, AMC23). These datasets span a broad complexity spectrum from basic arithmetic to competition-level mathematics, enabling a comprehensive evaluation of reasoning capabilities.
@@ 3.2 Baseline Results
Figure 2: Number of correct under 8 times rollout for different datasets.
We investigate how data difficulty influences the training dynamics of Qwen3 models. Specifically, we analyze the training convergence patterns through loss dynamics, accuracy trajectories, and generalization gaps, with three tiers of complexity (Easy, Medium, Hard). The detailed learning curves are shown in Figure 3.
@@
When focusing on the differences in learning efficiency between the unaligned Base model and the aligned model under the same experimental setting (as shown in Figure 3), the aligned models demonstrated a substantially higher initial accuracy and produced responses with a significantly longer average token length in the early stages of training. However, the performance improvement from additional learning steps of the aligned model was relatively modest, yielding only about a $2\%$ increase in accuracy. This result suggests that the current RL4LLM algorithm offers a slight improvement for aligned models that are already highly optimized.
@@ 4.1 Normalization
Advantage normalization is a well-established technique for reducing gradient variance and stabilizing policy optimization (Zheng et al., 2023), and it has become a standard component of RL training pipelines for language models. However, significant differences remain in how normalization is implemented. For example, GRPO (Shao et al., 2024) and RLOO (Ahmadian et al., 2024; Kool et al., 2019) use group-level normalization, calculating advantages relative to other responses within the same prompt to foster intra-context competition. On the other hand, REINFORCE++ (Hu et al., 2025) employs batch-level normalization, arguing that optimizing within a single prompt excessively can lead to reward hacking and hinder generalization, especially when response diversity is low.
@@
In contrast, batch normalization computes the reward over a rollout batch of size $N$ and $K$ sampled trajectories. The normalized advantage for the $i$-th response is:
@@
4.1.1 Advantage normalization is sensitive to reward mechanisms
Takeaway 1 Group-level normalization demonstrates robust efficiency across different reward settings. Batch-level normalisation provides more stable improvement under large scale reward setting.
To systematically evaluate the impact of advantage normalization on PPO variants with a value function using the Monte Carlo return target, we conducted experiments under a unified training framework, exploring three settings: no normalization, batch-level normalization, and group-level normalization. To highlight the differential impacts of the normalization techniques during training, we selected the Qwen3-base series models due to their low initial scores and high improvement potential (Yang et al., 2025). This choice ensures a fair comparison by minimizing confounding factors from alignment or prior optimization. Focusing on model scale as a key variable, we evaluate small (4B) and medium-sized (8B) models to empirically assess whether normalization techniques interact with model capacity. This approach allows us to derive practical insights into normalization strategies across different computational budgets and architectures.
[…]
4B-Base model with batch-level normalization Refer to caption 4B-Base model with group-level normalization Refer to caption
Figure 5: Top 2 rows: Accuracy over training iterations of Qwen3-4B-Base with batch-level normalization under different reward scale. The first row uses the easy training dataset, while the second row uses the medium training dataset. Bottom 2 rows: Accuracy over training iterations of Qwen3-4B-Base with group-level normalization under different reward scale.

### v2

@@
Abstract Reinforcement learning (RL) for LLM reasoning has rapidly emerged as a prominent research area, marked by a significant surge in related studies on both algorithmic innovations and practical applications. Despite this progress, several critical challenges remain, including the absence of standardized guidelines for applying RL techniques and a fragmented understanding of their underlying mechanisms. In addition, inconsistent experimental settings, variations in training data, and differences in model initialization have led to conflicting conclusions, obscuring the key characteristics of these techniques and creating confusion among practitioners when selecting appropriate techniques. This paper systematically reviews widely adopted RL techniques through rigorous reproductions and isolated evaluations within a unified open-source framework. We analyze the internal mechanisms, applicable scenarios, and core principles of each technique through fine-grained experiments, including datasets of varying difficulty, model sizes, and architectures. Based on these insights, we present clear guidelines for selecting RL techniques tailored to specific setups and provide a reliable roadmap for practitioners navigating the RL for the LLM domain. Finally, we show that a minimalist combination of two techniques can unlock the learning capability of critic-free policies with a vanilla PPO loss. The results demonstrate that our simple combination consistently improves performance, surpassing strategies such as GRPO and DAPO.
@@ 1 Introduction
Recent breakthroughs in large language models (LLMs) such as OpenAI o1 (Wu et al., 2024) and DeepSeek R1 (Shao et al., 2024) have positioned reinforcement learning (RL) as a key driver in unlocking advanced reasoning capabilities in LLMs. This is particularly evident in challenging reasoning tasks like mathematical problem solving (He et al., 2025a) and code generation (Zhuo et al., 2025), where RL has demonstrated the potential to elevate LLM performance beyond what pre-training alone can achieve. Such an emerging trend has sparked widespread interest within the research community in the direction of "RL for LLM" (or RL4LLM). In 2025, RL4LLM experienced explosive growth, producing hundreds of publications across arXiv and major conferences, covering a wide range of topics from algorithmic innovation to practical engineering solutions.
@@
However, this rapid progress is shadowed by the lack of usage guidelines for existing RL techniques and tricks (Huang et al., 2024a) as well as the absence of in-depth analysis of their underlying mechanisms. Specifically, these limitations can manifest as confusion among practitioners in choosing RL tricks, as different papers advocate conflicting solutions to the same problem. For instance, GRPO (Shao et al., 2024) advocates for group-level normalization to enhance policy stability, whereas REINFORCE++ (Hu et al., 2025) argues that batch-level normalization works better. Moreover, GRPO incorporates variance in normalization, yet Dr. GRPO (Liu et al., 2025a) explicitly recommends removing variance normalization to prevent bias. Similarly, GRPO (Shao et al., 2024) has achieved a breakthrough in performance through the strategy of using response-level loss calculation, while DAPO (Yu et al., 2025) has instead adopted token-level loss calculation. Such contradictory and chaotic phenomena underscore the fragmented understanding and inconsistent recommendations within the RL4LLM community. A likely cause for the above phenomenon is that the experimental settings, training data, and initialization of the existing work all have significant differences, which may also cause deviations in the summary of the conclusions.
@@
Apart from the confusion caused by the intrinsic differences of similar techniques, the numerous and seemingly orthogonal techniques, including Normalization, Clip, and Overlong Filtering, also increase the complexity of algorithm application in practice. Practitioners face non-trivial challenges in identifying an effective combination from a wide range of techniques to unlock the learning capacity of LLMs in specific scenarios. These ambiguities have naturally triggered a key requirement of practitioners:
@@
What scenarios are the existing techniques respectively suitable for? Is there a simple and generalizable combination that can be used to enhance policy optimization?
@@
Aligned with classic RL mechanism analysis methodologies (Andrychowicz et al., 2020; Engstrom et al., 2020; Huang et al., 2024a), we systematically review the widely used RL techniques by reproducing them and independently evaluating the actual impact of each technique, based on the same open-source infrastructure framework and policy models. To comprehensively cover practical scenarios, we design extensive experimental settings incorporating datasets of varying difficulty levels, diverse model sizes, and distinct model types. Furthermore, we conduct an in-depth analysis of their theoretical foundations, implementation details, and applicable scenarios as demons. The intuitive contribution is illustrated in Figure 1. Specifically, ❶ our empirical results reveal that most RL techniques exhibit strong preferences and sensitivities to the experimental setup, e.g., model type, data distribution, reward mechanism and hyperparameter. ❷ Based on the isolated analysis under our setup, we demonstrate that employing just two techniques, i.e., advantage normalization (group-level mean, batch-level std) and token-level loss aggregation, can unlock the learning capability of critic-free policies using vanilla PPO loss, surpassing mainstream RL4LLM algorithms incorporating redundant components. Our core contributions are selected as:
@@
1. Group-level normalization shows robust efficiency under each reward setting. Batch-level normalization provides more stable improvement under large scale reward setting. (§LABEL:sec.5.1.1) 2. Group-level mean and batch-level standard deviation enable further robust normalization. (§4.1.2) 3. Clip Higher promotes high-quality exploration for aligned models. (§4.2.1) 4. There appears to be a “scaling law” between the performance and the upper bound of the clipping on the small-sized model. (§4.2.3) 5. Compared to sequence-level loss aggregation, token-level aggregation is effective on base models but shows limited improvement on aligned models. (§4.3.1) 6. Overlong filtering enhances accuracy and clarity for short-to-medium reasoning tasks but provides limited benefits for long-tail reasoning. (§4.4.1) 7. Two techniques may unlock learning capacity in critic-free policies based on vanilla PPO loss. (§5)
@@ 2.2 Group Relative Policy Optimization (GRPO)
The effectiveness of the above normalization method can be understood from the perspective of reward shaping. By emphasizing the differences among candidate outputs for the same prompt, it effectively preserves the reliability of the gradient signal, even in sparse reward settings (Hu et al., 2020). Instead of adding a KL penalty to the reward, GRPO directly regularizes by directly adding the KL divergence between the trained policy and the reference policy to the loss. The overall surrogate objective is:
@@ 2.3 Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO)
Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025) is a recent RL method designed to address the unique challenges in LLM reasoning. For each question $q$ with gold answer $a$, DAPO samples a group of $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy, computes their rewards, and maximizes the following surrogate objective:
@@ 2.4 Reinforcement Learning Techniques
A variety of practical techniques have been introduced to stabilize optimization, reduce variance, and accelerate the convergence of LLMs on reasoning tasks. Drawing from prior research and practical implementations, we categorize commonly used techniques as follows.
@@
Clipping controls excessive updates in policy optimization and can be applied to rewards, advantages, or ratios. Furthermore, the Clip Ratio Higher (Yu et al., 2025) method relaxes the upper bound in PPO’s ratio clipping to better preserve exploration.
@@
Normalization of rewards or advantages helps stabilize gradient magnitudes. Representative approaches include: Batch-level Reward Normalization (Hu et al., 2025), Group-level Reward Normalization (Shao et al., 2024; Ahmadian et al., 2024), and Reward Shift without Standard Deviation (Liu et al., 2025a), which omits the standard deviation term to avoid difficulty bias.
@@
Filtering excludes uninformative or undesirable samples prior to gradient computation. Examples include: Overlong Filtering (Yu et al., 2025) to remove responses exceeding predefined length limits; Error Max Clip Mask and Right Min Clip Mask to filter overly incorrect or trivially correct samples; and Difficulty Mask (Yu et al., 2025; Zhang et al., 2025; Chu et al., 2025) to exclude samples outside a targeted difficulty range.
@@
The formulation of loss aggregation determines the relative contribution of each token to the overall objective. Common approaches include: Token-level Loss computes per-token advantages to reduce length bias, while Sequence-level Loss aggregates at the sequence level.
@@
These categories summarize the most prevalent strategies for improving RL in LLM reasoning. In this work, we focus on four key aspects: Normalization, Clipping, Masking, and Loss Aggregation, and conduct in-depth analyses of their mechanisms and practical utility.
@@ 3.1 Experimental Setup
We utilize the open-sourced ROLL framework [^1] Open source RL framework: https://github.com/alibaba/ROLL (Wang et al., 2025), an efficient and scalable platform specifically designed for reinforcement learning optimization in LLMs, to conduct all experiments. In addition, we adopt the PPO loss (Schulman et al., 2017), with advantage values computed using the REINFORCE algorithm (Sutton et al., 1999) as the unified RL baseline. To ensure consistency with prior research, we set the global batch size to $1024$ by using a rollout batch size of $128$ and sampling 88 responses per prompt, with a maximum response length of $8192$ tokens. The learning rate is set to 1e61e-6. For text generation, we use a top_p value of $0.99$, a top_k value of $100$, and a temperature of $0.99$.
@@
To comprehensively evaluate reinforcement learning (RL) techniques across parameter scales, our experiments cover two model sizes: Qwen3-4B and Qwen3-8B. For each model size, we include both non-aligned pre-trained versions (Qwen3-4B-Base and Qwen3-8B-Base) and aligned versions, enabling assessment RL gains from different initialization conditions [^2] Checkpoint links: https://huggingface.co/Qwen/Qwen3-4B; https://huggingface.co/Qwen/Qwen3-8B; https://huggingface.co/Qwen/Qwen3-4B-Base;https://huggingface.co/Qwen/Qwen3-8B-Base.
@@
To ensure reproducibility and fairness, we exclusively use open-source datasets for training, including SimpleRL-Zoo-Data (Zeng et al., 2025) and Deepmath (He et al., 2025a). To comprehensively examine how problem difficulty affects the RL technique’s performance, we randomly sample from the datasets, removing an excessive proportion of examples whose ground-truth label is simply “True” or “False”. This adjustment addresses the ostensible positive phenomenon, where models produce correct binary answers from erroneous reasoning chains, thereby introducing noisy supervision that compromises training quality (please refer to Appendix B.2 for case studies). Figure 2 visualizes the difficulty across the training dataset assessed by GPT-4o (Hurst et al., 2024).
@@
All the experiments are conducted on six math datasets: MATH-500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024), MinervaMath (Lewkowycz et al., 2022), and subsets of standardized examinations (AIME24-25, AMC23). These datasets span a broad complexity spectrum from basic arithmetic to competition-level mathematics, enabling a comprehensive evaluation of reasoning capabilities.
@@ 3.2 Baseline Results
Figure 2: Number of correct responses under 8 rollout iterations across datasets.
We investigate how data difficulty influences the training dynamics of Qwen3 models. Specifically, we analyze the training convergence patterns through loss dynamics, accuracy trajectories, and generalization gaps, across three tiers of complexity (Easy, Medium, Hard). The detailed learning curves are shown in Figure 3.
@@
When focusing on the differences in learning efficiency between the unaligned Base model and the aligned model under the same experimental setting (as shown in Figure 3), the aligned models exhibited substantially higher initial accuracy and produced responses with longer average token lengths during early training. However, additional learning yielded only modest gains, with accuracy improving by roughly $2\%$. This suggests that the current RL4LLM algorithm offers a slight improvement for aligned models that are already highly optimized.
@@ 4.1 Normalization
Advantage normalization is a well-established technique for reducing gradient variance and stabilizing policy optimization (Zheng et al., 2023), and it has become a standard component of RL training pipelines for language models. Yet, substantial differences remain in how it is implemented. For example, GRPO (Shao et al., 2024) and RLOO (Ahmadian et al., 2024; Kool et al., 2019) use group-level normalization, calculating advantages relative to other responses within the same prompt to foster intra-context competition. On the other hand, REINFORCE++ (Hu et al., 2025) employs batch-level normalization, arguing that optimizing within a single prompt excessively can lead to reward hacking and hinder generalization, especially when response diversity is low.
@@
In contrast, batch normalization computes the reward over a rollout batch of size NN with $K$ sampled trajectories. The normalized advantage for the $i$-th response is: